# Discovering Lie Groups with Flow Matching

### v1

@@ Abstract
Symmetry is fundamental to understanding physical systems, and at the same time, can improve performance and sample efficiency in machine learning. Both pursuits require knowledge of the underlying symmetries in data. To address this, we propose learning symmetries directly from data via flow matching on Lie groups. We formulate symmetry discovery as learning a distribution over a larger hypothesis group, such that the learned distribution matches the symmetries observed in data. Relative to previous works, our method, LieFlow , is more flexible in terms of the types of groups it can discover and requires fewer assumptions. Experiments on 2D and 3D point clouds demonstrate the successful discovery of discrete groups, including reflections by flow matching over the complex domain. We identify a key challenge where the symmetric arrangement of the target modes causes “last-minute convergence,” where samples remain stationary until relatively late in the flow, and introduce a novel interpolation scheme for flow matching for symmetry discovery.
@@ 1 Introduction
Symmetry has been central to mathematics and physics for over a century, and more recently, has featured prominently within machine learning. Classically, the underlying symmetry is first identified and categorized; Noether’s theorem (Noether1918) establishes the existence of corresponding conservation laws, which can be used to model and understand physical systems. In machine learning, symmetries serve as powerful inductive biases in equivariant neural networks (cohen2016group; kondor2018generalization; weiler2019e2cnn). These architectures leverage known symmetries in the data, such as rotational invariance in molecular structures (thomas2018tensor; satorras2021n) or translational equivariance in images (lecun1998gradient), to achieve superior performance with fewer parameters and training samples (bronstein2021geometric).
@@
A critical limitation in both physics and machine learning is that the exact symmetry group must be known a priori. In practice, the underlying symmetries are often unknown, approximate, or domain-specific. In physics, many systems exhibit hidden symmetries, from a simple 1D harmonic oscillator to black holes (gross1996role; liu2022machine). In material chemistry, non-obvious symmetries such as valency and electronegativity can reduce the combinatorial search space (davies2016computational). In computer vision, objects often contain local symmetries, and real-world scans are often partial or noisy (mitra2006partial; mitra2013symmetry). These applications share a common need: the ability to automatically discover and characterize symmetries from data, without relying on domain expertise or manual specification.
@@
Several prior works have focused on symmetry discovery in restricted settings, such as roto-translations in images (rao1998learning; miao2007learning), commutative Lie groups (cohen2014learning), or finite groups (zhou2020meta; karjol2024unified). Recent studies on learning Lie groups also have some limitations: benton2020learning assume a fixed Lie algebra basis and a uniform distribution over the coefficients; dehmamy2021automatic produce non-interpretable symmetries; yang2023generative assume a Gaussian distribution over the Lie algebra coefficients and use potentially unstable adversarial training; and allingham2024generative use a fixed Lie algebra basis and evaluate only over images.
@@
In this work, we propose LieFlow , which discovers symmetries via flow matching directly on Lie groups. We formulate symmetry discovery as learning a distribution over a larger symmetry group, which serves as a hypothesis space of symmetries, and the support of this distribution corresponds to the actual transformations observed in the data. Using flow matching allows us to capture both continuous and discrete symmetries within a unified framework and accurately model the highly multi-modal nature of symmetries. Unlike standard flow matching (lipman2023flow), which operates in data space, LieFlow learns flows directly over Lie group manifolds, conditioned on the data samples, enabling the generation of new plausible data that respect the underlying symmetry structure.
Our key contributions are to:
  • formulate symmetry discovery as a flow matching problem on Lie groups to match the true data symmetries,
  • provide a unified framework to discover continuous and discrete symmetries, including reflections via flow matching over the complex domain,
  • identify the “last-minute mode convergence” phenomenon, causing near-zero vector fields for most timesteps, and introducing a novel time schedule for symmetry discovery.
@@ 2 Preliminaries
A Lie group is a group that is also a smooth manifold, such that group multiplication and taking inverses are also differentiable. This makes them particularly suitable for describing continuous symmetries. For example, the group of planar rotations $\mathrm{SO}(2)$ is a Lie group consisting of one parameter $\theta$, where the rotation by angle $\theta$ can be represented as $R_{\theta}=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}$. Each Lie group $G$ is associated with a Lie algebra, denoted as $\mathfrak{g}$, which is the vector space tangent at the identity. The Lie algebra captures the local (infinitesimal) group structure and describes how the group behaves near the identity.
The exponential map $\exp:\mathfrak{g}\to G$ defines the relationship between the Lie algebra and the Lie group. For matrix Lie groups, which we consider exclusively in this work, the exponential map is the matrix exponential $\exp(A)=\sum_{k=0}^{\infty}\frac{A^{k}}{k!}$. If the exponential map is surjective, we can restrict its domain and define the logarithm map $\log:G\supset U\to\mathfrak{g}$ for the neighborhood $U$ around the identity such that $\exp(\log(g))=g,\forall g\in U$. Note that any Lie algebra element $A$ can be written as a linear combination of the Lie algebra basis, and the group element $\exp(A)$ is generated by exponentiating the linear combination. This allows us to generate the entire connected component of the Lie group using the Lie algebra, and thus the Lie algebra basis $L_{i}\in\mathfrak{g}$ is also called the (infinitesimal) generators of the Lie group.
@@
where $x_{t}$ and $\dot{x}_{t}$ come directly from the chosen interpolation.
This highlights the key advantage of FM: the transport dynamics can be learned entirely in a self-supervised manner, since interpolations between noise and data provide training targets without requiring access to likelihoods or score functions.
3 Related Work
Symmetry Discovery.
Many works on symmetry discovery differ on the types of symmetries that they can learn. Early works (rao1998learning; miao2007learning) used sequences of transformed images to learn Lie groups in an unsupervised way, but were limited to 2D roto-translations. Some recent works such as (zhou2020meta; karjol2024unified) can only discover finite groups, while others learn subsets of known groups (benton2020learning; romero2022learning; chatzipantazis2023learning). LieFlow is flexible enough to learn both continuous and discrete symmetries and highly multi-modal distributions. Other related works focus on continuous Lie groups, as in our work, but suffer from other limitations. dehmamy2021automatic uses Lie algebras in CNNs to discover continuous symmetries, but learn the symmetries end to end along with the task function, leading to non-interpretable symmetries. moskalev2022liegg extracts Lie group generators from a neural network to analyze the learned equivariance, which is different to our objective of learning which symmetries exist in the dataset. yang2023generative use GANs to learn the Lie algebra basis, but assume a known distribution, often Gaussian, over these elements. otto2023unified provide a unified framework based on the Lie derivative to discover symmetries of learned models via linear algebraic operators, whereas LieFlow focuses on learning distributions over group transformations from data. shaw2024symmetry propose a method based on vector fields that discovers continuous and discrete symmetries of machine learning functions beyond affine transformations, focusing on Killing type generators rather than probability flows on a Lie group as in LieFlow. ko2024learning learn infinitesimal generators of continuous symmetries from data via Neural ODEs and a task-specific validity score, capturing non-affine and approximate symmetries, whereas LieFlow models a full distribution over transformations within a prescribed matrix Lie group.
The closest to our method is allingham2024generative, where they first learn a canonicalization function in a self-supervised fashion and then learn a generative model that outputs the transformations for each prototype, given a large prior group. While the objective of learning a distribution over symmetries is similar, allingham2024generative is a two stage optimization process and uses costly maximum likelihood training. They also only evaluate on images, while we consider point clouds.
Generative Models on Manifolds.
Several recent works have extended CNFs to manifolds (gemici2016normalizing; mathieu2020riemannian; falorsi2021continuous) but rely on computationally expensive likelihood-based training. Some works (rozen2021moser; ben2022matching; chen2024flow) proposed simulation-free training methods, in particular, chen2024flow consider flow matching on general Riemannian manifolds. However, none of these works specifically consider Lie groups or the task of discovering symmetries from data. Other flow matching works directly incorporate symmetries (klein2023equivariant; song2023equivariant; bose2023se), but do not discover symmetry from data. Other works consider score matching on Lie groups (zhu2025trivialized; bertolini2025generative), but again assume a priori knowledge of the group. Closely related to our work is sherry2025flow, who define flow matching specifically for Lie groups. While methodologically similar to our work, they assume prior knowledge of the symmetry group and use specialized implementations for different Lie groups.
4 LieFlow: Symmetry Discovery via Flow Matching
@@ 4.1 Motivation and Formalism
A central challenge in symmetry discovery is to identify the subset of transformations actually expressed in data without assuming explicit prototypes or fixed augmentation rules. Existing methods often impose such structures externally, limiting scalability and generality. We address this by casting symmetry discovery as flow matching on Lie groups, which allows distributions over transformations to be learned directly from data while respecting the geometry of the underlying group.
@@
Our goal is to recover the unknown group of transformations $H$ which preserve the data distribution $q$ on data space $\mathcal{X}$. To discover the hidden group $H$, we first assume a large hypothesis group $G$ which acts on $\mathcal{X}$ that contains the most common symmetries. We wish to find the stabilizer subgroup $H\subset G$ such that $q(hx)=q(x)$ for all $x\in\mathcal{X}$ and $h\in H$.
@@
We frame the problem of finding $H$ as learning a distribution over the hypothesis group $G$ that concentrates on the subgroup $H$. To achieve this we train an FM model which flows a chosen prior distribution over $G$ to a distribution $p_{\theta}(g)$ supported on $H$. The structure of the discovered symmetries emerges from the concentration pattern of $p_{\theta}(g)$: continuous symmetries maintain a spread distribution across the manifold of admissible transformations, while discrete subgroups exhibit sharp peaks at a finite set of modes (e.g. the four rotations of $C_{4}$). Thus, LieFlow effectively filters the large hypothesis group down to the subgroup $H\subseteq G$ consistent with the data.
@@
We assume that the hypothesis group $G$ is connected, ensuring a well-defined logarithm map for compact groups. For the prior distribution $p$ over the group $G$, we use a uniform prior for compact groups and Gaussian priors over the Lie algebra coefficients for non-compact groups. When the exponential map is not surjective, we define the prior implicitly by sampling from the Lie algebra $\mathfrak{g}$ and applying the exponential map to construct group elements, ensuring well-defined training targets. In experiments, we consider $G=\mathrm{SO}(2)$ (planar rotations), $\mathrm{GL}(2)$ (invertible matrices) for the 2D datasets and $G=\mathrm{SO}(3)$ (3D rotations) for the 3D datasets. Although $\mathrm{SO}(d)$ and $\mathrm{GL}(d)$ for larger $d$ are certainly feasible in our framework, scaling our method to large or non-compact groups introduces additional numerical challenges, which we leave for future work.
Interpolant Paths Along Orbits.
Unlike either traditional FM models which operate in the Euclidean data space or Riemannian flows which operate entirely on $G$, we propose an FM model that is conditioned on data but outputs group transformations. More precisely, consider a data sample $x_{1}\sim q$ and a transformation $g\sim p(G)$ from the prior distribution. Let $x_{0}=gx_{1}$. We wish to learn to flow from $x_{0}$ to $x_{1}$ using only transformations in $G$. Let $A=\log(g^{-1})\in\mathfrak{g}$. We define the interpolation in the data space (an exponential curve) via the group action generated by $A$, where:
@@
This yields a probability path $p_{t}(x_{t}|x_{0})$ in data space that stays on the group orbit of $x_{1}$.
@@
The derivative of the path $\frac{d}{dt}x_{t}\in T_{x_{t}}\mathcal{X}$ defines a vector field along $x_{t}$. This defines a flow over $\mathcal{X}$ which stays within the orbit of $x_{1}$, only moving in directions dictated by the group action. To achieve symmetry discovery, we must learn both the flow from $x_{0}$ to $x_{1}$ and reconstruct the transformation $g$ that generated it. For convenience, we configure the flow network to output the Lie algebra element $A$ instead of $\frac{d}{dt}x_{t}$. We can then update $x_{t}$ by scaling and exponentiating $A$ and then applying the group action as in Algorithm 2. This allows us to reconstruct $g=\exp(-A)$. That is, we set $u_{t}(x_{t}\ |x_{1})=A$. (Note that $A$ pushes forward to the velocity vector $\frac{d}{dt}x_{t}$ under the group action. See Appendix A for a more detailed derivation.)
@@
$\mathcal{L}_{\mathrm{LieCFM}}(\theta)=\mathbb{E}_{t,x_{1}\sim q,x_{t}\sim p_{t}(x_{t}\mid x_{1})}\|v_{t}^{\theta}(x_{t})-u_{t}(x_{t}\mid x_{1})\|^{2}_{\mathcal{G}},$ (5)
where $x_{t}$ follows the exponential curve defined previously, $p_{t}(x_{t}\mid x_{1})$ is the distribution of points along the exponential curves, and $\mathcal{G}$ is a Riemannian metric on the Lie group, which equips the manifold with an inner product on each tangent space, allowing us to measure distances along the manifold. We use a left-invariant metric as it is completely determined by the inner product on the Lie algebra $\mathfrak{g}$ and the push-forward of the left action is an isometry, making computations simpler. The specific form of this metric is determined by the structure of the Lie group under consideration.

### v2

# arXiv:2512.20043v2[cs.AI] 01 Mar 2026

@@ Abstract
Symmetry is fundamental to understanding physical systems and can improve performance and sample efficiency in machine learning. Both pursuits require knowledge of the underlying symmetries in data, yet discovering these symmetries automatically is challenging. We propose LieFlow, a novel framework that reframes symmetry discovery as a distribution learning problem on Lie groups. Instead of searching for the symmetry generators, our approach operates directly in group space, modeling a symmetry distribution over a large hypothesis group $G$. The support of the learned distribution reveals the underlying symmetry group $H\subseteq G$. Unlike previous works, LieFlow can discover both continuous and discrete symmetries within a unified framework, without assuming a fixed Lie algebra basis or a specific distribution over the group elements. Experiments on synthetic 2D and 3D point clouds and ModelNet10 show that LieFlow accurately discovers continuous and discrete subgroups, significantly outperforming a state-of-the-art baseline, LieGAN, in identifying discrete symmetries.
Machine Learning, ICML
@@ 1 Introduction
Symmetry has long been central to mathematics and physics and, more recently, has featured prominently within machine learning. Classically, to model physical systems, the underlying symmetry is first identified and categorized; Noether’s theorem then (Noether, 1918) establishes the existence of corresponding conservation laws, which can be used to understand and predict the system. In machine learning, symmetries serve as powerful inductive biases in equivariant neural networks (Cohen and Welling, 2016; Kondor and Trivedi, 2018; Weiler and Cesa, 2019). These architectures leverage known symmetries in the data, such as rotational invariance in molecular structures (Thomas et al., 2018; Satorras et al., 2021) or translational equivariance in images (LeCun et al., 1998), to achieve superior performance with fewer parameters and training samples (Bronstein et al., 2021).
Figure 1: LieFlow discovers symmetries from data by assuming a hypothesis Lie group $G$ and learning the distribution of transformations over $G$ via flow matching over the Lie group, conditioned on data samples $x$.
@@
A critical limitation in both physics and machine learning is that the exact symmetry group must be known a priori. In practice, the underlying symmetries are often unknown, approximate, or domain-specific, e.g. hidden symmetries in physical systems (Gross, 1996; Liu and Tegmark, 2022), non-obvious chemical properties (Davies et al., 2016), and partial symmetries in 3D geometry (Mitra et al., 2006, 2013). Many applications can thus benefit from the ability to automatically discover and characterize symmetries from data, without relying on domain expertise or manual specification.
@@
Several prior works have focused on symmetry discovery in restricted settings, such as roto-translations in images (Rao and Ruderman, 1998; Miao and Rao, 2007), commutative Lie groups (Cohen and Welling, 2014), or finite groups (Zhou et al., 2020; Karjol et al., 2024). Recent studies on learning Lie groups also have some limitations: Benton et al. (2020) assume a fixed Lie algebra basis and a uniform distribution over the coefficients; Dehmamy et al. (2021) produce non-interpretable symmetries; Yang et al. (2023) assume a Gaussian distribution over the Lie algebra coefficients which is largely limited to continuous groups and use potentially unstable adversarial training; and Allingham et al. (2024) use a fixed Lie algebra basis and evaluate only over images.
@@
In this work, we propose LieFlow, a new framework that fundamentally reframes symmetry discovery as a distribution learning problem on Lie groups. Conceptually, this shifts the paradigm of estimating generators to recovering group structure via support learning in group space. Specifically, we learn a distribution over a larger symmetry group, which serves as a hypothesis space of symmetries. The support of this distribution corresponds to the actual transformations observed in the data. We achieve this via flow matching directly on Lie groups (Figure 1).
Unlike standard flow matching (Lipman et al., 2023), which operates in data space, LieFlow learns a flow on a Lie group $G$ which is acting on the data space, conditioned on a data point, enabling the generation of new plausible data that respect the underlying symmetry structure. Thanks to the flexibility of flow matching, LieFlow can capture both continuous and discrete symmetries within a unified framework and accurately model the multi-modal nature of distributions supported on subgroups. We observe that discrete symmetry groups can be particularly challenging since cancellation among the different flow vectors results in small velocity until late in the flow, creating a “last-minute convergence” phenomenon. We counter this issue by proposing a novel time schedule for flow matching.
Qualitative and quantitative experimental results on synthetic 2D and 3D datasets and real-world dataset ModelNet10 (Wu et al., 2015) show that LieFlow can discover discrete and continuous symmetries accurately, and achieve significant improvements over LieGAN (Yang et al., 2023) baseline in identifying discrete symmetries.
Our key contributions are as follows:
  • We propose a novel formulation of symmetry discovery as a flow matching problem on Lie groups. Our formulation covers continuous, discrete, approximate, and partially observed symmetries.
  • We identify a “last-minute convergence” phenomenon for discrete symmetries and propose a novel time scheduling method to address it.
  • Extensive experiments on synthetic and real-world datasets demonstrate that LieFlow can accurately discover underlying symmetry groups, and achieve significant improvements over LieGAN (Yang et al., 2023) baseline in identifying discrete symmetries.
2 Related Work
Symmetry Discovery.
Many works on symmetry discovery differ on the types of symmetries that they can learn. Early works (Rao and Ruderman, 1998; Miao and Rao, 2007) used sequences of transformed images to learn Lie groups in an unsupervised way, but were limited to 2D roto-translations. Some recent works (Zhou et al., 2020; Karjol et al., 2024) can only discover finite groups, while others learn partially observed symmetries, i.e. subsets of known groups (Benton et al., 2020; Romero and Lohit, 2022; Chatzipantazis and Pertigkiozoglou, 2023). Shaw et al. (2024) discover continuous and discrete symmetries of machine learning functions beyond affine transformations, but focus on Killing type generators rather than probability flows on a Lie group. In contrast, LieFlow is flexible enough to learn both continuous and discrete symmetries, highly multi-modal distributions, and also partially observed symmetries via flow matching.
Several related works focus on continuous Lie groups but suffer from other limitations. Dehmamy et al. (2021) use Lie algebras in CNNs but learn the symmetries end-to-end along with the task function, leading to non-interpretable symmetries. Moskalev et al. (2022) extract Lie group generators from a neural network to analyze the learned equivariance, rather than discovering which symmetries exist in the dataset. Otto et al. (2023) discover symmetries of learned models via linear algebraic operators. Ko et al. (2024) learn infinitesimal generators of continuous symmetries from data via Neural ODEs and a task-specific validity score, capturing non-affine and approximate symmetries. Unlike these approaches, LieFlow models a full distribution over transformations within a prescribed matrix Lie group.
Yang et al. (2023) and Allingham et al. (2024) are most related to our work as they also learn distributions over transformations. Yang et al. (2023) propose LieGAN, which uses adversarial training to learn the Lie algebra basis while assuming a fixed distribution over the coefficients, typically Gaussian. This makes LieGAN well-suited for continuous symmetries; while it can discover discrete groups, doing so requires a carefully chosen prior distribution such as a multi-modal delta mixture. Allingham et al. (2024) first learn a canonicalization function to obtain a prototype as the orbit representative, then learn a generative model over transformations conditioned on the orbit representative. LieFlow, on the other hand, learns a marginal transformation distribution across all orbits directly via flow matching, without assuming a fixed distribution or learning explicit prototypes.
Generative Models on Manifolds.
Several recent works have extended continuous normalizing flows to manifolds (Gemici et al., 2016; Mathieu and Nickel, 2020; Falorsi, 2021) but rely on computationally expensive likelihood-based training. Some works (Rozen et al., 2021; Ben-Hamu et al., 2022; Chen and Lipman, 2024) have proposed simulation-free training methods; in particular, Chen and Lipman (2024) considers flow matching on general Riemannian manifolds. However, none of these works specifically consider Lie groups or the task of discovering symmetries from data. Other flow matching works directly incorporate known symmetries (Klein et al., 2023; Song et al., 2023; Bose et al., 2023) and do not infer symmetry from data. Another class of methods considers score matching and flow matching directly on Lie groups as data spaces (Zhu et al., 2025; Bertolini et al., 2025; Sherry and Smets, 2025). While methodologically similar to our work, they assume prior knowledge of the symmetry group and use specialized implementations for different Lie groups.
@@ 3 Preliminaries
A Lie group is a group that is also a smooth manifold, such that group multiplication and inversion are smooth maps. For example, the group of planar rotations $\mathrm{SO}(2)$ is a Lie group parameterized by angle $\theta$, with elements $R_{\theta}=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}$. Each Lie group $G$ has an associated Lie algebra $\mathfrak{g}$, which is the tangent space at the identity. The Lie algebra captures the local (infinitesimal) group structure.
The exponential map $\exp\colon\mathfrak{g}\to G$ maps the Lie algebra to the Lie group. For matrix Lie groups, which we consider exclusively in this work, the exponential map is the matrix exponential $\exp(A)=\sum_{k=0}^{\infty}\frac{A^{k}}{k!}$. If the exponential map is surjective, we can restrict its domain and define the logarithm map $\log\colon G\supset U\to\mathfrak{g}$ for the neighborhood $U$ around the identity such that $\exp(\log(g))=g$ for $g\in U$. Note that any Lie algebra element $A$ can be written as a linear combination of the Lie algebra basis. This allows us to generate the entire connected component of the identity in the Lie group using the Lie algebra, and thus the Lie algebra basis $L_{i}\in\mathfrak{g}$ is also called the (infinitesimal) generators of the Lie group.
@@
where $x_{t}$ and $\dot{x}_{t}$ come from the chosen interpolation.
A key advantage of FM is that the transport dynamics can be learned entirely in a self-supervised manner, since interpolations between noise and data provide training targets without requiring access to likelihoods or score functions.
4 Symmetry Discovery via Flow Matching
@@ 4.1 Motivation and Formalism
A central challenge in symmetry discovery is not merely to estimate transformation parameters, but to determine which transformations are symmetries at all. In realistic settings, we are not given paired observations of original and transformed data, and the space of candidate transformations is typically much larger than the true symmetry group expressed in the data. We address this challenge by reframing symmetry discovery as a distribution learning problem on groups. Rather than directly estimating generators or invariants, we learn a probability distribution over a large hypothesis group $G$, whose support concentrates on the true symmetry subgroup $H\subseteq G$. We realize this idea using flow matching on Lie groups, which provides a flexible and self-supervised mechanism for learning which transformations preserve the data distribution.
@@
Our goal is to recover the unknown group of transformations $H$ that preserves the data distribution $q$ on data space $\mathcal{X}$. We first assume a large hypothesis group $G$ acting on $\mathcal{X}$ that contains hypothetical symmetries. We wish to find the subgroup $H\subseteq G$ such that $q(hx)=q(x)$ for all $x\in\mathcal{X}$ and $h\in H$.
@@
We frame the problem of finding $H$ as learning a distribution over the hypothesis group $G$ that concentrates on the subgroup $H$. To achieve this, we train an FM model that maps a chosen prior distribution over $G$ to a distribution $p_{\theta}(g)$ supported on $H$. The structure of the discovered symmetries emerges from the concentration pattern of $p_{\theta}(g)$: continuous symmetries yield a distribution spreading across the manifold of valid transformations, while discrete subgroups exhibit sharp peaks at a finite set of group elements (e.g. the four-fold rotations of $C_{4}$). Thus, LieFlow effectively filters the large hypothesis group down to the subgroup $H\subseteq G$ consistent with the data.
@@
We assume that the hypothesis group $G$ is connected, ensuring a well-defined logarithm map for compact groups. For the prior distribution $p$ over the group $G$, we use a uniform prior for compact groups and bounded priors over the Lie algebra coefficients for non-compact groups. When the exponential map is not surjective, we define the prior implicitly by sampling from the Lie algebra $\mathfrak{g}$ and applying the exponential map to construct group elements, ensuring well-defined training targets. In experiments, we consider $G=\mathrm{SO}(2)$ (planar rotations), $\mathrm{GL}(2)$ (invertible matrices) for the 2D datasets and $G=\mathrm{SO}(3)$ (3D rotations) for the 3D datasets.
Interpolation Paths Along Orbits.
We now describe how to construct interpolation paths that respect the group structure. Consider a data sample $x_{1}\sim q$ and a transformation $g\sim p(G)$ drawn from the prior distribution. We apply the transformation to obtain $x_{0}=gx_{1}$, which serves as the starting point of our flow. We wish to learn to flow from $x_{0}$ to $x_{1}$ using only transformations in $G$. Let $A=\log(g^{-1})\in\mathfrak{g}$ be the Lie algebra element that generates the inverse transformation. We define the interpolation in the data space via the group action:
@@
At $t=0$, we have $x_{0}=gx_{1}$, and at $t=1$, we recover $x_{1}$ since $\exp(A)=g^{-1}$. Marginalizing over $g$, this defines a conditional $p_{t}(x_{t}\mid x_{1})$ that stays on the group orbit of $x_{1}$.
@@
The derivative $\frac{d}{dt}x_{t}\in T_{x_{t}}\mathcal{X}$ defines a vector field along the path $\{x_{t}\}_{t\in[0,1]}$. Unlike standard flow matching, where the network outputs a velocity in data space, we configure our network to output a Lie algebra element $A\in\mathfrak{g}$ directly. That is, we set the conditional target field to $u_{t}(x_{t}\ |x_{1})=A$.
This choice has two advantages. First, the update rule $x_{t+\Delta_{t}}=\exp(\Delta_{t}A_{t})x_{t}$ ensures that the trajectory stays on the group orbit of $x_{1}$. Second, we can accumulate the transformations to obtain group elements in $H$ (Line 9, Algorithm 2). Note that $A$ is constant along the path, so the network learns to predict the Lie algebra element generating the transformation from $x_{0}$ to $x_{1}$, regardless of $t$. The velocity field $\frac{d}{dt}x_{t}$ is obtained from $A$ via the differential of the group action; see Appendix A for more details.
@@
$\mathcal{L}_{\mathrm{LieCFM}}=\mathbb{E}_{t,q,p_{t}(x_{t}\mid x_{1})}\|v_{t}^{\theta}(x_{t})-u_{t}(x_{t}\mid x_{1})\|^{2}_{\mathcal{G}},$ (5)
where $x_{1}\sim q$, $x_{t}$ is defined by Eq. (4), $p_{t}(x_{t}\mid x_{1})$ is the distribution of points along the paths, and $\mathcal{G}$ is a Riemannian metric on the Lie group, which equips the manifold with an inner product on each tangent space, allowing us to measure distances along the manifold. We use a left-invariant metric as it is completely determined by the inner product on the Lie algebra $\mathfrak{g}$ and the push-forward of the left action is an isometry, making computations simpler. The specific form of this metric is determined by the structure of the Lie group under consideration.
Although transformations are sampled from a broad hypothesis group during training, the flow-matching objective implicitly filters out transformations that are not true symmetries of the data. When a sampled transformation preserves the data distribution, applying it to a data point produces another statistically valid sample, and the resulting interpolation paths induce consistent training signals across many data points. In contrast, transformations that do not preserve the data distribution produce inconsistent targets: different data points yield incompatible directions, which cancel out when the model averages training signals. As a result, only transformations that are globally consistent with the data admit a coherent vector field, and the learned symmetry distribution concentrates its probability mass on the true symmetry group.